Calculate Reynolds Number for flow of water at a speed of 1 cm/s in a circular channel with a diameter of
100
Answer #1

Reynolds number is 1. the flow is laminar as reynolds number is less than 2000.

e s

Where ? is the density of the fluid,
V is the velocity of the fluid,
? is the density of fluid,
? is the viscosity of fluid,
L is the length or diameter of the fluid.

take density as 998kg/m3 and viscosity as .001 kg/m.s.
Answer #2

Re = DV/v = (100*10^-6 * 0.01)/(10^-6) = 1,

turbulent flow